Calculate current through individual resistors in the following circuit diagram50 Ω 20 Ω 30 Ω + 12V 12 13 15 Ω 20 Ω 30 V (14) i4 15 - 80 Ω 20 V 60 Ω
Answer #1

If you like my answer plz give thumbs up.

Loop1:

-30+50i1+20(i1-i2)+80(i1-i4)=0

Loop2:

20(i2-i1)+30(i2-i3)+15(i2-i4)=0

Loop3:

12+20(i3-i5)+30(i3-i2)=0

Loop4:

-20+80(i4-i1)+15(i4-i2)=0

Loop5:

20+20(i5-i3)+60i5=0

From loop2 and loop4:

If we solve we get equation in terms of i3 ,i2 and i5.if we substitute in loop3 we get equation in terms of only i3 and i2,similarly solve loop 3 and loop 5 we get in terms of i1 and i4 ao fibally of we solve those equations we get answer.